OpenStudy (anonymous):

Oil spilled from a tanker spreads in a circle whose circumference increases at a rate of 40ft/sec. How fast is the area of the spill increasing when the circumference of the cirlce is 100pi ft

OpenStudy (amistre64):

when dealing with rates of change, we wanna ask, what can we derive

OpenStudy (anonymous):

ok

hero (hero):

Why is it everytime I find a question no one is working on, someone shows up....This person has been waiting for 13 minutes...no response. I show up, and within a minute amistre shows up..... :S

OpenStudy (amistre64):

dC/dt = 40 dA/dt = dA/dr * dr/dC * dC/dt ... maybe

OpenStudy (amistre64):

im a ninja lol

hero (hero):

I'm Mafia

OpenStudy (anonymous):

everyone help me

OpenStudy (amistre64):

C = 2pi r C/2pi = r A = pi (C/2pi)^2 dA/dt = pi 2(C/2pi) (1/2pi) dC/dt

OpenStudy (amistre64):

\[\frac{dA}{dt} = \frac{C(40)}{2pi}\] \[\frac{dA}{dt} = \frac{C(20)}{pi}\]and C = 100 right?

OpenStudy (amistre64):

100pi

OpenStudy (anonymous):

100pi

OpenStudy (amistre64):

so pis cancels and we got 100(20) = 2000 units per time

OpenStudy (anonymous):

... huh

OpenStudy (amistre64):

huh isnt something i can relate to; your gonna have to be a bit more specific

OpenStudy (anonymous):

lol sry can u write out the stops using the equation tab

OpenStudy (anonymous):

C=40FT How fast is the area incraesing when C=100pi ft?

OpenStudy (amistre64):

\[C = 2pi\ r\] \[\frac{C}{2pi} = r\] \[A = pi \left(\frac{C}{2pi}\right)^2\] \[A = pi \frac{C^2}{4pi\ pi}\] \[A = \frac{C^2}{4pi}\] now to derive

OpenStudy (anonymous):

thanks

OpenStudy (amistre64):

\[\frac{d}{dt}\left(A = \frac{C^2}{4pi}\right)\] \[\frac{dA}{dt} = \frac{2C}{4pi}\ \frac{dC}{dt}\] \[\frac{dA}{dt} = \frac{C}{2pi}\ \frac{dC}{dt}\] C=100pi and dC/dt = 40 \[\frac{dA}{dt} = \frac{100pi(40)}{2pi}\] \[\frac{dA}{dt} = 50(40)=2000\]

OpenStudy (amistre64):

and this is in feet/sec

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you sooo much i will review this

OpenStudy (anonymous):

A two line Mathematica solution with comments is attached. Confirms amistre64's result.